What volume (in milliliters) of a 0.635 M NaOH solution contains 8.35 g of NaOH?
1
Expert's answer
2020-11-09T13:53:27-0500

"\begin{aligned}
&\textsf{mass of }NaOH = 8.35g\\
&\textsf{amount of }NaOH = \frac{8.35}{40} = 0.209mol\\
&\textsf{concentration of }NaOH = 0.635M
\end{aligned}"


"Volume = \dfrac{no.\ of\ moles}{concentration}"


"\begin{aligned}
Volume =\dfrac{0.209}{0.635}= 0.329dm^3
\end{aligned}"


Volume = "0.329dm^3" = 0.329L = "0.329\u00d7 1000" mL = 329mL


"\therefore" The volume (in milliliters) of the solution is 329 mL.